# Reconfigurable Intelligent Surfaces for 6G Mobile Networks: An Industry R&D Perspective

### v2

@@ Abstract
The reconfigurable intelligent surface (RIS) technology is a potential solution to enhance network capacity and coverage without significant investment into additional infrastructure in 6G networks. This work highlights the interest of the mobile communication industry on RIS, and discusses the development of liquid crystal-based RIS for improved energy efficiency and coverage in the millimeter-wave band. Furthermore, the paper discusses perspectives and insights from an industry R&D point of view, addressing relevant use cases, technical requirements, implementation challenges, and practical considerations for RIS deployment optimization in the context of 6G networks.
@@
6G, reconfigurable intelligent surfaces, network, liquid crystal
@@ I Introduction
The fifth-generation (5G) mobile networks are currently being rolled out worldwide. Theoretically, 5G should provide data rates of up to 10 Gbps in average and latencies below 10 ms. However, it is clear that with future applications, for example in connection with telepresence, mixed realities or for highly specialized use in an industrial context, the re-quirements for data rate, reliability, connection density and latency will continue to increase. Among other targets, 6G aims at larger transmission rates, lower latencies, and more energy efficiency. It provides connectivity even under unfavorable conditions in order to enable safety-critical applications such as the connection for unmanned autonomous vehicles. These targets cannot be achieved by simply deploying additional base stations, but requires alternative, innovative solutions. With regard to the long-term national and international climate goals as well as the social acceptance of 6G, it is crucial to focus on energy efficiency in all necessary developments and network expansions. Sustainability is highlighted as an essential criterion in which 6G should build upon 5G.
@@
Potential solutions to achieve 6G’s targets include utilizing new spectrum and implementing further network densification with distributed MIMO concepts [1]. However, these measures result in increased energy requirements, which can only be mitigated by developing more energy-efficient components. Transitioning to higher frequencies also leads to increased propagation losses, requiring extremely dense deployments with direct line of sight (LOS). Additionally, existing networks may not easily support future applications with different radio coverage requirements, such as aerial connectivity, necessitating network coverage extension to the lower airspace. Cooperative relaying systems [2] have also been proposed as a potential solution to improve coverage and have even been incorporated in LTE-Advanced [3]. However, they require reliable backhaul connection, which increases its implementation complexity and cost.
@@
A novel approach to enhance network capacity, coverage, and range without additional base stations or repeaters is through reconfigurable intelligent surfaces (RIS) [4, 5]. RIS involves using passive controllable surfaces to shape the radio environment, influencing wave propagation for various purposes such as improving cell coverage and controlling interference. RIS can also enhance spatial diversity and channel capacity in MIMO systems, and find potential applications to emerging technologies, such as Internet of Things, Industry 4.0, ground-to-air communications, localization, sensing, and physical layer security [4].
@@
Because of its potential to significantly improve the performance of radio access network, RISs have caught the attention of academia, industry, funding bodies, and regulators, and, therefore, significant investments have been directed towards the development and evaluation of this technology towards 6G. For example, the 6G LICRIS (Liquid Crystal Reconfigurable Intelligent Surfaces) project aims to develop RIS based on liquid crystal (LC) for surface reconfigurability from a holistic network perspective. Compared to semiconductor-based RIS, LC is more energy efficient in higher frequencies, it is scalable, cost-effective, and a promising technology candidate for RIS hardware implementation.
@@
The objective of this article is to share the authors’ perspective and learnings on RIS coming from an industry R&D point of view. More specifically, relevant use cases for future investigation, their technical requirements, and existing challenges are discussed in Section II. Similar considerations are presented in [6, 7], but the present work differentiates itself from the existing literature by diving deeper into implementation and deployment aspects. In particular, the hardware implementation of novel LC-based concepts are presented in Section III. The practical problem of channel modeling, RIS deployment optimization and dimensioning are discussed in Section IV. These problems are discussed along novel insights into the line-of-sight coverage improvement of RIS in urban deplyoment and the corresponding RIS size optimization. Finally, the paper is concluded with potential topics for further investigation that shall support the incorporation of RIS in mobile communication standards.
@@ II-A Relevant Use Cases
Use cases for RIS are currently discussed in the radio communication industry, starting from coverage extension towards smart antenna applications. However, standardisation bodies, such as 3GPP or ETSI have just started to analyse possible opportunities. One of the first standard developing organizations (SDO) working group to look into RIS specifications and applications has been established in ETSI early in 2021, namely the ETSI ISG RIS (ETSI industry specification group for RIS). Still, the technical group reports are not published and in a drafting stage. So far, they have identified the following potential use cases for RIS deployments:
@@
    Signal enhancement for blind users: the direct LOS is obstructed by a building. This reduces the signal quality on receiver side massively. By using the nearby installed RIS, a “quasi-LOS” connection can be established and the signal quality is significantly improved. This application is particularly important in high-frequency scenarios (mmWave and THz communication), as these kind of signals are highly susceptible to blockage and pathloss [8].
@@
    Interference suppression: By deploying RISs on cell edges the desired signal strengths can be improved compared to the inter-cell interference (interference by other neighboring base stations/cells) signal strengths. This is done by passive beamforming as the signals are fed to the target direction only. Following this the spread of signals into areas that should not actually be covered is reduced [9].
@@
    Improve channel rank conditions: RISs allow to manipulate the wireless channel by adding additional propagation paths into the desired direction. By this the channel rank condition can be improved [9].
@@
    Physical layer security: In addition to the above mentioned use cases, RISs can be used to improve the secrecy capacity of communication networks [10]. The secrecy capacity is a common metric in physical layer security. It is defined as the difference between the link capacity of the legitimate transmitter/receiver and the link capacity of the eavesdropper transmitter/receiver. These link capacities are calculated based on the corresponding signal-to-noise (SNR) ratios. Due to the beam shaping capability of an RIS, the SNR for the legitimate link can be improved and the SNR for eavesdropper is reduced [11].
@@
    Mounting of RIS on unmanned aerial vehicle (UAV) To assist and enhance terrestrial communication, RISs can be attached to UAVs [12]. This idea is similar to the mounting of RIS modules to buildings. However, attachment to UAVs provides even more degrees of freedom with regard to positioning. It extends the coverage of the wireless signals by beamforming with UAV trajectory optimization [9].
@@
    RIS to support simultaneous wireless information and power transfer (SWIPT): Internet of things (IoT) devices like small sensors that support our todays life are more and more spread into the world. Traditionally these devices are equipped with small batteries. One emerging approach studies the power supply of these devices by using the radiated energy for the wireless information transfer (SWIPT). By introducing RISs in such scenarios the energy collecting and data transmission capabilities can be boosted [9].
@@
    Enhancement of localization and sensing capabilities by using RIS: By introducing RIS into localization and sensing applications, new synchronization location refer-ences and configurations are generated. Furthermore, new geometric measurements can be done and the localization accuracy is improved [13].
@@
    RIS assisted non-orthogonal multiple access (NOMA): Compared to conventional multiple input multiple output (MIMO)-NOMA, RIS assisted NOMA can improve the throughout, spectral efficiency, and user connectivity by overcoming the challenges of random channel fluctuations, blockage and user mobility [14].
@@
    Enhancing indoor signal strength: Due to the intrinsic impedance difference between free-space and glass, the signal magnitude through windows is reduced drastically. To overcome this challenge, transparent reflecting intelligent surfaces, also known as transparent intelligent surfaces, represent a suitable solution for improving the signal quality without greatly impairing the actual function of the glass, leading to an optical transparency of 80% [15].
Other applications are expected to follow along with the development of meta-materials allowing RIS design and embedding RIS products into the ecosystem of radio communication systems by standardisation bodies.
@@ II-B1 Regulation
The European baseline for the regulation of radio equipment sold and deployed in the European marketplace, including Germany in particular, is the European Radio Equipment Directive 2014/53/EU, commonly known as “RED” [16].
@@
As RIS modules and RIS based products will be primarily deployed as radio network components to extend coverage, as depicted in Figure 1, or offer radio beam shaping capabilities to transmitters, they obviously have to meet regulatory requirements according to RED. Thus, it is expected that there will be the need for RIS specific harmonized requirements standardization in the European Union. Still, there is no request by the European Commission, however, standardization bod-ies like ETSI have already started into this, namely the ETSI ISG RIS.
@@
Among other stakeholders, Rohde & Schwarz is a contributing member of this ETSI ISG RIS group to identify potential regulatory testing needs at the earliest possible stage. The common knowledge about regulatory testing will be contributed to this project in order to specify potential testing requirements and adjust all RIS testing investigations accordingly. A very early outcome of the discussion on RIS testing needs in the industry has been the common conclusion towards primarily over-the-air testing in anechoic environments known from legacy antenna test systems. Compared to antennas, where the radiated signal is typically fed via antenna connectors, the “feed signal” for an RIS module is the impinging radio signal. Consequently, the entire measurement setup is radiated (feed and probe signal) or “over-the-air”. Therefore, legacy antenna measurement systems need to be extended by feed signal antennas in addition to the probe antennas, see Figure 2.
Regulatory requirements basically focus on the protection of other radio services and the most effecient usage of the given spectrum as an important asset. Thus, it is expected to see regulatory requirements including unwanted and spurious emissions caused by the RIS deployment as well as blocking and intermodulation rejection requirements.
@@ II-B2 Conformance
Conformance testing is regarded as one of the key success factors in mobile communication in the last 3 decades. The idea is a common industry driven regime to test and certify mobile devices and network components, for instance, according to the common standard, provided by 3GPP (3rd Generation Partnership project). Conformance testing goes far beyond regulatory testing, as its focus is on ensuring seamless operation of the devices worldwide, since radio regulations may differ in different geographical regions. As RIS might become an important component of future radio communication network architectures, RIS products will have to join that conformance testing regime. Thus, beside regulatory assessments, most probably there will be additional, performance-oriented and standard-compliance oriented requirements and testing needs.
As soon as RIS will be used to manipulate the impinging signal with respect to radio parameters, e.g., beamforming, and signal quality parameters,e.g., modulation quality, there is a need to test those capabilities according to the standard. So far, as explained before, SDOs are still in an early phase regarding RIS specifications, and the potential need for conformance and performance assessment is yet to be specified.
There are various ways to look at RIS deployments: For instance, RIS as an intelligent reflector to direct the impinging signal to a wanted, other than specular, reflection angle. Thus, tests to assess that capability are to be expected. Another view could be regarding the RIS as an intelligent antenna, manipulating the impinging wave with respect to radio beam characteristics or its modulation quality. Consequently, tests assessing the radio beam or signal quality would be required. Independent on the way an RIS is used, it seems obvious that all conformance and performance tests are over-the-air (OTA) tests, just as all expected regulatory assessments. The important question, however, is whether there is a need to specify new, RIS-specific assessment metrics. As the RIS manipulates an impinging signal in a certain way, by default, there might be a need not only to qualify its performance by looking at the reflected signal only, but taking the comparison with the impinging wave into account.
@@ II-B3 Operational Requirements Gaps
Possibly the most serious concern of the industry regarding the operation of RIS is the fact that it will not be irradiated by wanted signals only, but also any other radio signal out there. Thus, co-existence issues need to be addressed with great interest. According to regulations, radio equipment shall not harm other services, thus, co-existence can be regarded as a regulatory issue, too. Therefore, coexistence testing scenarios need to be considered, including in-band (wanted signal spectrum) and out of band, e.g., adjacent spectrum signals. Nevertheless, this is obviously a wide area and needs to be addressed by dedicated investigations, which is out of the scope of this work.

### v4

@@ Abstract
The reconfigurable intelligent surface (RIS) technology is a potential solution to enhance network capacity and coverage without significant investment in additional infrastructure in 6G networks. This work highlights the interest of the mobile communication industry in RIS, and discusses the development of liquid crystal-based RIS for improved energy efficiency and coverage in the millimeter-wave band. Furthermore, the paper discusses perspectives and insights from an industry R&D point of view, addressing relevant use cases, technical requirements, implementation challenges, and practical considerations for RIS deployment optimization in the context of 6G networks. A hardware design of an RIS with liquid crystal at 28 GHz is presented. A propagation model for RIS as a new part of the system architecture is discussed, with approaches of semi-empirical models, geometric models, and their combination through the application of artificial intelligence/machine learning. Finally, a channel model for deployment optimization and dimensioning is presented, with the findings that a rather large RIS is favorable for coverage improvement, as well as greater attenuation at higher frequencies combined with a smaller RIS size.
@@
6G, reconfigurable intelligent surfaces, network, liquid crystal, unit cell, network densification, coverage issues, channel models, propagation modeling, over-the-air (OTA) -testing.
publicationid: pubid: DOI: 10.1109/ACCESS.2024.3485227 ©2024 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission. See http://www.ieee.org/publications standards/publications/rights/index.html for more information.
@@ I Introduction
The fifth-generation (5G) mobile networks are currently being rolled out worldwide. Theoretically, 5G should provide data rates of up to 10 Gbps on average and latencies below 10 ms. However, it is clear that with future applications, for example in connection with telepresence, mixed realities, or for highly specialized use in an industrial context, the requirements for data rate, reliability, connection density, and latency will continue to increase. Among other targets, the sixth-generation (6G) aims at larger transmission rates, lower latencies, and more energy efficiency. It provides connectivity even under unfavorable conditions to enable safety-critical applications such as the connection for unmanned autonomous vehicles. These targets cannot be achieved by simply deploying additional base stations, but require alternative, innovative solutions. With regard to the long-term national and international climate goals, as well as the social acceptance of 6G, it is crucial to focus on energy efficiency in all necessary developments and network expansions. Sustainability is highlighted as an essential criterion on which 6G should build upon in comparison to 5G.
@@
Potential solutions to achieve 6G’s targets include utilizing new spectrum and implementing further network densification with distributed Multiple-Input Multiple-Output (MIMO) systems concepts [1]. However, these measures result in an increase in energy requirements, which can only be mitigated by developing more energy-efficient components. Transitioning to higher frequencies also leads to increased propagation losses, requiring extremely dense deployments with direct line of sight (LOS). Additionally, existing networks may not easily support future applications with different radio coverage requirements, such as aerial connectivity, necessitating network coverage extension to the lower airspace. Cooperative relaying systems [2] have also been proposed as a potential solution to improve coverage and have even been incorporated in Long-term Evolution (LTE) Advanced [3]. However, they require reliable backhaul connection, which increases its implementation complexity and cost.
@@
A novel approach to enhance network capacity, coverage, and range without additional base stations or repeaters is through reconfigurable intelligent surfaces (RIS) [4, 5]. RIS involves using passive controllable surfaces to shape the radio environment, influencing wave propagation for various purposes such as improving cell coverage and controlling interference. RIS can also enhance spatial diversity and channel capacity in MIMO systems and find potential applications in emerging technologies such as the Internet of Things (IoT), Industry 4.0, ground-to-air communications, localization, sensing, and physical layer security[4]. Due to its potential to significantly improve the performance of radio access networks, RISs have caught the attention of academia, industry, funding bodies, and regulators. Consequently, significant investments have been directed towards the development and evaluation of this technology for 6G. Industry players such as ZTE [6, 7, 8], AGC [9], SK Telecom [10], and NTT [11] have started to invest in the research and development of RIS to expand their product portfolios.
@@
For example, the 6G-LICRIS (Liquid Crystal Reconfigurable Intelligent Surfaces) project aims to develop RIS based on liquid crystal (LC) for surface reconfigurability from a holistic network perspective. Compared to semiconductor-based RIS, LC is more energy efficient in higher frequencies, scalable, cost-effective, and a promising technology candidate for RIS hardware implementation. The aim of the 6G-LICRIS project is to develop reconfigurable intelligent surfaces and integrate them into the mobile network. To achieve this, new materials based on liquid crystals are being developed. The newly developed technology will be integrated into a test mobile network to study radio wave propagation using the RIS. Finally, the integration of RIS into a real radio environment to demonstrate end-to-end data transmission in the actual mobile network is considered [12].
The project has identified several key steps to improve its features and categorizations. It will follow the European Telecommunications Standards Institute (ETSI) reports and use metamaterial and reflectarray structures. The RIS will have a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas. It will operate in the mmWave frequency range with time division duplex (TDD) duplex mode. The RIS’s goal is to enhance coverage, spectral efficiency, and beam management while improving energy efficiency. The RIS’s deployment scenario can be static in both indoor and outdoor environments. It can have a centralized or distributed management system with a network-controlled mode of operation. The communication and networking topology can be both indoor and outdoor, and the RIS should focus on improving indoor signal strength from an outdoor base station. While the RIS will not have sensing or localization functionalities, it should make use of channel estimation and precoding offboard. Key performance indicators should include rate or throughput, coverage, and energy efficiency, which depend on design parameters like size, carrier frequency, and bandwidth. The connectivity and reliability through deployment scenarios should be improved to utilize single or multiple RISs controlled individually to maximize coverage.
@@
The objective of this article is to share the authors’ perspective and learnings on RIS coming from an industry research and development (R&D) point of view at the beginning of the 6G-LICRIS project. More specifically, relevant use cases for future investigation, their technical requirements, and existing challenges are discussed in Section II. Similar considerations are presented in [13, 14], but the present work differentiates itself from the existing literature by diving deeper into implementation and deployment aspects. In particular, the hardware implementation of novel LC-based concepts are presented in Section III. The practical problem of channel modeling, RIS deployment optimization and dimensioning are discussed in Section IV. These problems are discussed with novel insights into the line-of-sight coverage improvement of RIS in urban deployment and the corresponding RIS size optimization. Finally, the paper is concluded with potential topics for further investigation that shall support the incorporation of RIS in mobile communication standards.
@@ II-A Relevant Use Cases
Use cases for RIS are currently being discussed in the radio communication industry, starting from coverage extension towards smart antenna applications. However, standardisation bodies, such as 3rd Generation Partnership Project (3GPP) or ETSI have just started to analyze possible opportunities. One of the first standard developing organizations (SDO) working group to look into RIS specifications and applications has been established in ETSI early in 2021, namely the ETSI ISG RIS (ETSI Industry Specification Group for RIS). Still, the technical group reports are not published and in a drafting stage. So far, they have identified the following potential use cases for RIS deployments:
@@
    Signal enhancement for hidden users: The direct LOS is obstructed by a building. This reduces the signal quality on receiver side massively. By using the nearby installed RIS, a “quasi-LOS” connection can be established, and the signal quality is significantly improved. This application is particularly important in high-frequency scenarios (mmWave and THz communication), as these kind of signals are highly susceptible to blockage and pathloss [15].
@@
    Interference suppression: By deploying RISs on cell edges, the desired signal strengths can be improved compared to the inter-cell interference (interference by other neighboring base stations/cells) signal strengths. This is done by passive beamforming, as the signals are fed to the target direction only. Following this the spread of signals into areas that should not actually be covered is reduced [16].
@@
    Increase network capacity: The quality of signal in networks with time-varying channels can be improved [17] as well as in wide-band applications [18]. For improving the channel rank conditions, RISs allow manipulation of the wireless channel by adding additional propagation paths into the desired direction. By this the channel rank can be improved [16].
@@
    Physical layer security: In addition to the above mentioned use cases, RISs can be used to improve the secrecy capacity of communication networks [19]. The secrecy capacity is a common metric in physical layer security. It is defined as the difference between the link capacity of the legitimate transmitter/receiver and the link capacity of the eavesdropper transmitter/receiver. These link capacities are calculated based on the corresponding signal-to-noise ratios (SNR). Due to the beam shaping capability of an RIS, the SNR for the legitimate link can be improved and the SNR for eavesdropper is reduced [20].
@@
    Mounting of RIS on unmanned aerial vehicle (UAVs): To assist and enhance terrestrial communication, RISs can be attached to UAVs [21]. This idea is similar to the mounting of RIS modules to buildings. However, attaching them to UAVs provides even more degrees of freedom with regard to positioning. It extends the coverage of the wireless signals by utilizing beamforming with UAV trajectory optimization [16].
@@
    RIS to support simultaneous wireless information and power transfer (SWIPT): Internet of Things (IoT) devices like small sensors that support our daily life, are increasingly deployed worldwide. Traditionally, these devices are equipped with small batteries. One emerging approach studies the power supply of these devices by using the radiated energy for the wireless information transfer (SWIPT). By introducing RISs in such scenarios, the energy harvesting and data transmission capabilities can be boosted [16, 22, 23].
@@
    Enhancement of localization and sensing capabilities using RIS: The introduction of RIS into localization and sensing applications generates new synchronization location references and configurations. Additionally, it enables new geometric measurements and improves localization accuracy [24].
@@
    RIS-assisted Non-Orthogonal Multiple Access (NOMA): Compared to conventional Multiple Input Multiple Output (MIMO)-NOMA, RIS-assisted NOMA can improve throughput, spectral efficiency, and user connectivity by overcoming challenges such as random channel fluctuations, blockages, and user mobility [25].
@@
    Enhancing indoor signal strength: Due to the intrinsic impedance difference between free-space and glass, the signal magnitude through windows is drastically reduced. To overcome this challenge, transparent reflecting intelligent surfaces, also known as transparent intelligent surfaces, provide a suitable solution for improving signal quality without greatly impairing the actual function of the glass, achieving an optical transparency of 80% [26].
Other applications are expected to follow, with developments in metamaterials enabling RIS design and the integration of RIS products into the ecosystem of radio communication systems by standardization bodies.
@@ II-B1 Regulation
The baseline for regulating radio equipment sold and deployed in the European marketplace, including Germany, is the European Radio Equipment Directive 2014/53/EU, commonly known as “RED” [27].
@@
As RIS modules and RIS-based products will primarily be deployed as radio network components to extend coverage, as depicted in Figure 1, or offer radio beam shaping capabilities to transmitters, they must meet the regulatory requirements according to RED. Therefore, it is expected that there will be a need for RIS-specific harmonized requirements standardization in the European Union. Although there is no request from the European Commission yet, standardization bodies like ETSI have already started this process, namely the ETSI ISG RIS.
@@
Among other stakeholders, Rohde & Schwarz is a contributing member of the ETSI ISG RIS group, aiming to identify potential regulatory testing needs at the earliest possible stage. The common knowledge about regulatory testing will be contributed to the 6G-LICRIS project to specify potential testing requirements and adjust all RIS testing investigations accordingly. A very early outcome of the discussion on RIS testing needs in the industry has been the common conclusion toward primarily over-the-air testing in anechoic environments, similar to those known from legacy antenna test systems. Compared to antennas, where the radiated signal is typically fed via antenna connectors, the “feed signal” for an RIS module is the impinging radio signal. Consequently, the entire measurement setup is radiated (feed and probe signal) or “over-the-air”. Therefore, legacy antenna measurement systems need to be extended by feed signal antennas in addition to the probe antennas, as depicted in Figure 2.
Regulatory requirements primarily focus on the protection of other radio services and the most efficient usage of the given spectrum as an important asset. Thus, it is expected that regulatory requirements will include addressing unwanted and spurious emissions caused by RIS deployment, as well as blocking and intermodulation rejection requirements.
@@ II-B2 Conformance
Conformance testing is regarded as one of the key success factors in mobile communication over the last three decades. The idea is a common industry-driven regime to test and certify mobile devices and network components, for instance, according to the common standard provided by 3GPP. Conformance testing goes far beyond regulatory testing, as its focus is on ensuring the seamless operation of devices worldwide, since radio regulations may differ across different geographical regions. As RIS might become an important component of future radio communication network architectures, RIS products will need to adhere to that conformance testing regime. Therefore, in addition to regulatory assessments, there will most likely be additional, performance-oriented and standard-compliance-oriented requirements and testing needs.
As soon as RIS is used to manipulate the impinging signal with respect to radio parameters, such as beamforming, and signal quality parameters, such as modulation quality, there will be a need to test those capabilities according to the standard. So far, as previously explained, SDOs are still in an early phase regarding RIS specifications, and the potential need for conformance and performance assessment has yet to be specified.
There are various ways to look at RIS deployments: for instance, an RIS can operate as an intelligent reflector to direct the impinging signal to a desired, non-specular reflection angle. Therefore, tests to assess this capability are to be expected. Another perspective could consider the RIS as an intelligent antenna, manipulating the impinging wave with respect to radio beam characteristics or its modulation quality. Consequently, tests assessing the radio beam or signal quality would be required. Regardless of how an RIS is utilized, it seems clear that all conformance and performance tests, as well as all expected regulatory assessments, should be OTA tests. The important question, however, is whether there is a need to specify new, RIS-specific assessment metrics. As the RIS manipulates an impinging signal in a specific way, by default, there might be a need not only to qualify its performance by examining the reflected signal alone but also by considering the comparison with the impinging wave.
@@ II-B3 Operational Requirements Gaps
Possibly the most serious concern of the industry regarding the operation of RIS is the fact that it will be exposed not only to desired signals but also to any other radio signal out there. Thus, co-existence issues need to be addressed with great interest. According to regulations, radio equipment shall not interfere with other services, therefore, co-existence can also be regarded as a regulatory issue. In view of this, coexistence testing scenarios need to be considered, including both in-band (desired signal spectrum) and out-of-band, such as adjacent spectrum signals. Nevertheless, this is a broad area and needs to be addressed by dedicated investigations, which are beyond the scope of this work.